OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 4 − (x + 2) = 6 − 5(x + 4) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

First get rid of parenthesis. Use distributive property.

OpenStudy (anonymous):

Step 1. Get rid of the parenthasis. The equation would look like this 4 - x - 2 = 6 - 5x - 20 Step 2. Simplify the arithatic. The equation would look like this 2 - x = -14 - 5x Step 3. Subtract all terms from the left side that we don't want to solve for from both sides. (everything but x) The equation would look like this 2 - x - 2 = -14 - 5x - 2 Step 4. Simplify the arithmatic again. The equation would look like this -x = -16 - 5x Step 5. Get common terms we are solving for to the side we designated by subtracting from both sides. (bring all x terms to the left side) The equation would look like this -x - (- 5x) = -16 - 5x - (-5x) Note that a - (-5x) = +5x Step 6. Simplify arithmatic. The equation would look like this 4x = -16 Step 7. Divide by the constant infront of the variable to solve for the variable by itself. (divide both sides by 4) The equation would look like this -4x/(4) = -16/(4) Step 8. Simplify arithmatic. The equation would look like this x = -16/4 or simplified as x = -4

OpenStudy (anonymous):

thanks great help !!